40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 28 04:27:33 2016, ericq, Update, General, mucking about alscheck.pdfweird_olpit.pdf
    Reply  Thu Sep 29 09:49:53 2016, rana, Update, General, mucking about 
    Reply  Sat Oct 1 10:03:28 2016, ericq, Update, General, mucking about 
       Reply  Tue Oct 4 02:59:48 2016, ericq, Update, General, mucking about 2016-10-04_ALSXspectra.pdf
Message ID: 12527     Entry time: Sat Oct 1 10:03:28 2016     In reply to: 12521     Reply to this: 12529
Author: ericq 
Type: Update 
Category: General 
Subject: mucking about 

Some things I did last night:

I measured the X PDH OLG, and turned the gain down by ~6dB to bring the UGF back to 10kHz, ~50deg phase margin, 10dB gain margin. However, the error signal on the oscilloscope remained pretty ratty. Zooming in, it was dominated by glitches occuring at 120Hz. I went to hook up the SR785 to the control signal monitor to see what the spectrum of these glitches looked like, but weirdly enough connecting the SR785's input made the glitches go away. In fact, with one end of a BNC connector plugged into a floating SR785 input, touching the other end's shield to any of the BNC shields on the uPDH chassis made the glitches go away.

This suggested some ground loop shenanigans to me; everything in the little green PDH shelves is plugged into a power strip which is itself plugged into a power strip at the X end electronics rack, behind all of the sorensens. I tried plugging the power strip into some different places (including over by the chamber where the laser and green refl PD are powered), but nothing made the glitches go away. In fact, it often resulted in being unable to lock the PDH loop for unknown reasons. This remains unsolved.

As Gautam and Johannes observed, the X green beat was puny. By hooking up a fast scope directly to the beat PD output, I was able to fine tune the alignment to get a 80mVpp beat, which I think is substaintially bigger than what we used to have. (Is this plus the PDH gain changed really attributable to arm loss reduction? Hm)

However, the DFD I and Q outputs have intermittent glitches that are big enough to saturate the ADC when the whitening filters are on, even with 0dB whitening gain, which makes it hard to see any real ALS noise above a few tens of Hz or so. Turning off the whitening and cranking up the whitening gain still shows a reasonably elevated spectrum from the glitches. (I left a DTT instance with a spectrum on in on the desktop, but forgot to export...) The glitches are not uniformly spaced at 120Hz as in the PDH error signal. However, the transmitted green power also showed intermittant quick drops. This also remains unsolved for the time being. 

ELOG V3.1.3-